OpenStudy (anonymous):

Find the probability of the following five-card poker hands from a 52-card deck. In poker, aces are either high or low. Two pair (2 cards of one value, 2 of another value)

OpenStudy (valpey):

First you need the number of ways to have two pair, then you need the number of ways to have a five card hand. For the number of ways to have two pair you have: 13 possible first pairs * 6 possible suit combinations * 12 possible second pairs * 6 possible suit combinations /2 because JJ44 is the same as having 44JJ * 11 ranks which aren't one of your two pair * 4 suits So that's 13*12*11*6*6*4/2 for the number of ways to get two pair or \[\frac{13*12*11*6*6*4}{2} = {13*12*11*6*6*2}\] For the number of ways to have a 5 card hand is 52 choose 5 or \[\frac{52!}{5!*47!} = \frac{52*51*50*49*48}{5*4*3*2} = {52*51*5*49*4}\] That gives us a probability of \[\frac{13*12*11*6*6*2}{52*51*5*49*4}\]

OpenStudy (anonymous):

hmm for some reason i just dont get it? my answers to choose from are all fractions.. im not getting fractions with my answer?

OpenStudy (valpey):

Convert the answer we are working with into a decimal, then convert your possible answers into decimals and it will work out.

OpenStudy (valpey):

But forget the multiple choice part for a second. Lets talk about the number of ways to have a five card hand. Do you see where that comes from?

OpenStudy (anonymous):

no i dont): i dont get this question AT ALL!

OpenStudy (anonymous):

would the answer by chance happen to be C.3432/433115?

OpenStudy (valpey):

@brittkay12, I'll feel a lot better about helping if we can get the understanding down first, then talk about the answer. Lets imagine that you have a deck of just seven cards, and they are lettered A,B,C,D,E,F,G. How many ways can I make a three-card hand with those seven cards?

OpenStudy (anonymous):

twice?

OpenStudy (valpey):

Some examples of three card hands might be: ABC, ADG, DAB, DGE, CAB, DBC

OpenStudy (valpey):

It is okay if you aren't sure, I will help you through. Just let me know what you are thinking.

OpenStudy (anonymous):

oh.. i dont know then): my answer choices are A.20592/433115 B.22464/433115 C.3432/433115 D.5148/433115 im thinking its either C or A.. but it coud also be B or D.): i dont have enough time left for all this tho.): am i better off guessing A, or C?!

OpenStudy (valpey):

Please don't guess. I promise this explanation won't take as long as it seems like it will.

OpenStudy (valpey):

Back to the three card hands. You might think of it like: Well there are seven possible cards I could pick for the first card, six possible cards left to draw for the second card and 5 possible cards I could draw for the third card.

OpenStudy (anonymous):

i believe ima just guess C. im on a time limit & running out of time & i still have more questions to answer..!

OpenStudy (valpey):

Good luck with your studies.

OpenStudy (anonymous):

thankyou for trying!!! although a simple answer would have been appreciated!

OpenStudy (valpey):

This isn't a site for people to do your homework for you.

OpenStudy (anonymous):

goodbye.